OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 5x + 1.

OpenStudy (bakonloverk):

just solve these y = 5 * -2 + 1 y = 5 * 0 + 1 y = 5 * 2 + 1

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."